Circular No. 7912

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002dd Z. Tsvetanov, J. Blakeslee, and H. Ford, Johns Hopkins University; D.

Magee and G. Illingworth, University of California, Santa Cruz; A. Riess, Space Telescope Science Institute; and the ACS Science Team report the discovery of a second apparent supernova (cf. SN 2002dc, IAUC 7908) in observations of the Hubble Deep Field North taken with the Advanced Camera for Surveys on the Hubble Space Telescope. SN 2002dd is located at α = 12^{h}36^{m}55^{s}.35, δ = +62^{o}12′46′′.1 (equinox 2000.0). The new object was detected in exposures taken with the F775W filter (at m_{AB} = 24.00 ± 0.12) on May 11.39 UT, with the G800L grism on May 11.52, and with the F850LP filter (at m_{AB} = 24.15 ± 0.12) on May 21.17; the on-orbit photometric calibration of the ACS in these bandpasses has not yet been completed, and the quoted errors are dominated by the zero-point uncertainty, estimated at 0.1 mag. The object was not present in WFPC2 F814W images of the HDFN taken with the HST in Dec. 1995. The spectrum extracted from the ACS G800L grism exposure appears to be that of a type-Ia supernova at a redshift z = 1.06 observed near maximum. The spectrum shows pronounced absorption features due to Ca II H and K and the ironpeak elements (Fe, Co). SN 2002dd is located 1′′.6 east and 0′′.6 north of the nearest bright reference galaxy (listed as object 3-486.0 in Williams et al. 1996, A.J. 112, 1335); however, this galaxy cannot be the host, as it has a measured redshift of z = 0.79 (Cohen et al. 1996, Ap.J. 471, L5). No clear host galaxy can be identified at present.

GX 339–4 D. M. Smith, University of California, Berkeley, on behalf of a collab-

oration (including T. Belloni, W. A. Heindl, E. Kalemci, R. Remillard, M. Nowak, J. H. Swank, and S. Corbel), reports that the blackhole candidate GX 339–4, in outburst since Mar. 26, entered the rare Very High State (VHS) of x-ray emission around May 6. The energy spectrum is dominated by a disk blackbody but has a strong power-law tail. The power spectrum of the fast variability switches among at least three states: band-limited noise with a broad quasiperiodic oscillation (QPO) near 9 Hz, red noise with a narrower QPO near 6 Hz, and a quiet state with almost no variability. RXTE/PCA observations are taking place at least twice weekly. The RXTE ASM shows that the count rate has started to decline from its peak. Monitoring in all other wavebands is strongly encouraged; the VHS was last observed in 1988 (Miyamoto et al. 1991, Ap.J. 383, 784).

2002 May 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green